OpenStudy (anonymous):

Solve for the unknown: |2x-5|x-2|| = |x+6| + 3

OpenStudy (anonymous):

what a pain

OpenStudy (anonymous):

let me see if i can work this out on pencil and paper before writing the solution.

OpenStudy (anonymous):

sure. We know it's a pain, even with 4 brains already working on it. Thanks for the help.

OpenStudy (anonymous):

i think you have to work in cases. for example the right hand side is a piecewise function, namely \[ |x+6|+3 = \left\{ \begin{array}{lr} x+9 & : x >6\\ -x-3 & : x < 6 \end{array} \right.\]

OpenStudy (anonymous):

left hand side is more of a pain

OpenStudy (anonymous):

ok i have one of the solutions.

OpenStudy (anonymous):

shoot. :)

OpenStudy (anonymous):

the left hand side is |2x-5|x-2||

OpenStudy (anonymous):

now if x > 2 this is |2x-5(x-2)| = |2x-5x+10| = |-3x+10| yes?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

because of course if x > 2 then |x-2|=x-2

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

now |-3x+10| = -3x + 10 if x < 10/3 and |-3x+10|=10-3x if x > 10/3

OpenStudy (anonymous):

so if x> 6 both conditions hold and we have the equation \[10-3x=x+9\] solve for x to get \[x=\frac{19}{2}\]

OpenStudy (anonymous):

that is if x > 6 then the left hand side is 10-3x and the right hand side is x+9 so solve for x and that is one answer. now we have to do it again for the other side

OpenStudy (anonymous):

go on. :)

OpenStudy (anonymous):

again on the left if x < 2 then |2x-5|x-2|| is |2x-5(2-x)| is |2x-10+5x| is |7x-10|

OpenStudy (anonymous):

hope these steps are clear

OpenStudy (anonymous):

i am just replacing |x-2| with (2-x) for x < 2

OpenStudy (anonymous):

yes yes. go on

OpenStudy (anonymous):

and |7x-10| = 7x-10 if x > 10/7 and |7x-10|=10-7x if x < 10/7

OpenStudy (anonymous):

now for x < 10/7 still x > -6 so set \[10-7x=x+9\] \[1=8x\] \[x=\frac{1}{8}\]

OpenStudy (anonymous):

and those are the two solutions

OpenStudy (anonymous):

what a pain!

OpenStudy (anonymous):

btw i made a typo on the first line

OpenStudy (anonymous):

yes we got that

OpenStudy (anonymous):

*bows* :))) thanks man, that was your second medal :D

OpenStudy (anonymous):

i was so impressed with myself for writing a piecewise function that i wrote it incorrectly. it should be \[|x+6|+3 = \left\{ \begin{array}{lr} x+9 & : x >-6\\ -x-3 & : x < -6 \end{array} \right.\]

OpenStudy (anonymous):

i had 6 instead of -6

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i guess while i am showing off that you should say the line \[7x-10\] for x between 10/7 and 2 does not intersect the line x+9

OpenStudy (anonymous):

anyway, we got the main gist of your solution, and was able to work it out. thanks again :-bd